Chemistry (3 part Questions)

3a. Use the following data table and equation to generate a standard curve for the concentration of the product [Fe(SCN)]+2 by plotting equilibrium concentration versus absorbance. A graph is required for credit including title, labeled axis, and correct linear regression line(s) ?

Fe+3 + SCN-1 <--> [Fe(SCN)]+2

Test Tube 0.100 M FeCl3 0.00100M KSCN Water Absorbance
1 10.00 mL 1.00 mL 9.00 mL 0.012
2 10.00 2.00 8.00 0.101
3 10.00 3.00 7.00 0.245
4 10.00 4.00 6.00 0.338
5 10.00 5.00 5.00 0.456
6 10.00 6.00 4.00 0.557
7 10.00 7.00 3.00 0.793
8 10.00 8.00 2.00 0.848

3b.Calculate the equilibrium concentration of a sample of [Fe(SCN)]+2 with an absorbance of 0.194 ?

3c. Calculate the equilibrium constant for a reaction with an initial concentration of 0.178 M Fe+3 and 0.153 M SCN-1 with an absorbance at the equilibrium of 0.194 ?

I have no idea how to do this problem, the first part needs a graph (Step by step on how u label and graph it) and the 2nd and 3rd part Idk how to do it. Please show a graph and do a step by step on how to solve for the answer. Your help is always greatly appreciated! Thank you in advance.
Answer #1

For the equillibrium reaction:

Fe3+ + SCN- [Fe(SCN)]2+

let us assume that whole of the limiting reagent is consumed during the reaction. Since product is a colored compound, will show absorbance in visivle range. From the given concentrations it is apparent that SCN- solution has much lower concentration in comparison to Fe3+. therefore it is assumed that whole of SCN- will react and moreover the stoichiometric ratio of SCN- to product is 1:1, hence same concentration of product will form .i.e initial [SCN-] = equillibrium concentration of [Fe(SCN)]2+

Step 1: Calculate the concentration of reactants in the reaction mixture as:

M1*V1 = M2*V2

where M1 nad M2 are the concentration of stock solution used and concentration in reaction mixture, respectively

V1 and V2 are the volumes of stock solution used and final (total) volume of the reaction mixture = 20 mL (from the table provided)

example or test tube 1:

0.1*10=M2*20

or [Fe3+] = 0.05M

Similarly fot [SCN-]:

0.001*1=M2*20

or [SCN-] = 5*10-5 M (therefore its the limiting reagent in the reaction)

See the table below for all the test solutions and corresponding plot.

3 b. Calculate the equilibrium concentration of a sample of [Fe(SCN)]+2 with an absorbance of 0.194 ?

Using the Lambert-Beer's Law as:

A = Cl

where A=absorbance, = absorptivity coefficient, C = concentartion of [Fe(SCN)]+2 , l=path length (=1 cm)

From the above plot of Absorbance vs concentration and fitting it to a linear function as y=mx+c where m = slope = el = 20000 and

c = intercept = 0. Therefore is calculated as:

* 1 cm=20000 M-1

= 20000 M-1cm-1

Now using the same Lambert-Beer's Law and A= 0.194, = 20000 M-1cm-1 and l= 1 cm, C is calculated as:

0.194 = 20000 *C*1

or C= 9.7 *10-6 M

Therefore, the equilibrium concentration of a sample of [Fe(SCN)]+2 with an absorbance of 0.194 is 9.7 * 10-6 M

3c. Calculate equilibrium constant for a reaction with an initial concentration of 0.178 M Fe+3 and 0.153 M SCN-1 with an absorbance at the equilibrium of 0.194.

Ans: For the reaction equillibrium constant , keq = [Fe(SCN)]+2/[Fe+3]*[SCN-1]

where concentrations of reactants and product are at equillibrium

from part b the equilibrium concentration of a sample of [Fe(SCN)]+2 with an absorbance of 0.194 is 9.7 * 10-6 M.

Here the concentration of reactants are comparable thus, there equillibrium concentartions are calculated as follows:

From absorbance values, [Fe(SCN)]+2 = 9.7 * 10-6 M =x

therefore, at equillibrium [Fe+3] = 0.178-x = 0.178 - 9.7 * 10-6 M

and [SCN-1] = 0.153-x = 0.153 - 9.7 * 10-6 M

substituting in above formula of equillibrium constant we get:

keq = 9.7 * 10-6/[(0.178 - 9.7 * 10-6)*(0.178 - 9.7 * 10-6)]

or keq= 3.56 * 10-4

Therefore, equilibrium constant for a reaction with an initial concentration of 0.178 M Fe+3 and 0.153 M SCN-1 with an absorbance at the equilibrium of 0.194 is 3.56 * 10-4 .